Parth (parthkohli):

\[\int {\sin^5 x dx} \]

Parth (parthkohli):

So, should I start with the following?\[\int \left(\sin(x)\right)^4 \sin x dx\]

Parth (parthkohli):

I think I might be able to use substitution rule in this one.

OpenStudy (anonymous):

yes so u must create some cos from somewhere

Parth (parthkohli):

Oh... wait... yes!\[\int (\sin^2(x))^2\sin xdx \]Now to use the identity\[\sin^2 x = 1 - \cos^2x \]

Parth (parthkohli):

Right?

OpenStudy (anonymous):

exactly

Parth (parthkohli):

\[\int (1 - \cos ^2x)^2 - \sin x dx \]

Parth (parthkohli):

Hmm... \(u = \cos x \) would work?

mathslover (mathslover):

Yes ... let u = cos x

OpenStudy (anonymous):

yes

Parth (parthkohli):

Let's check.\[du = \cos x dx\]

mathslover (mathslover):

Oops sorry for interrupting

Parth (parthkohli):

Sorry, sinxdx ^^

Parth (parthkohli):

Yeah!! Works!!

Parth (parthkohli):

\[\int (1 - u^2)^2 du \]Right?

Parth (parthkohli):

Oh, then to simplify (1 - u^2)^2?

OpenStudy (anonymous):

sorry to butt in; but you can also use the reduction formulae; much faster.

Parth (parthkohli):

Haven't heard of it, Mishi.

Parth (parthkohli):

\[\int 1 - 2u^2 + u^4 du \]

Parth (parthkohli):

lol, now I got it :)

Parth (parthkohli):

Whom should I thank?

mathslover (mathslover):

Thank mukushla ...

Parth (parthkohli):

\[u - {2 \over 3}u^3 + {1 \over 5}u^5+c\]

Parth (parthkohli):

OK, now to substitute back for u... I should be able to do that... thank you :D